OpenStudy (natsent06):

Ned traveled 1/4 the distance that Ben traveled. If Ned traveled 12 miles, how far did Ben travel? A) 3 miles B) 16 miles C) 48 miles D) 24 miles

OpenStudy (anonymous):

c) 48 miles

OpenStudy (texaschic101):

Ned traveled 1/4 the distance of Ben N = (1/4)B Ned traveled 12 miles so we sub 12 = (1/4) B -- divide by 1/4 12/(1/4) = B 12 * 4 = B 48 = B check.. N = (1/4)B 12 = 1/4(48) 12 = 48/4 12 = 12 (correct) Ben traveled 48 miles